Two charges, a= +6uC and qz=-5uC.

are separated by a distance r = 15cm.

What is the magnitude of the electric

force? Is the force repulsive or attractive?


1
Expert's answer
2022-02-28T13:40:50-0500

The Coulomb's law says


"F=k\frac{q_1q_2}{r^2}\\
=9*10^9*\frac{6*10^{-6}*5*10^{-6}}{0.15^2}=12\;\rm N"